Can you help me with this equation problem?
A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 100 gal of solution of 2% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 5% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?

Answer by ankor@dixie-net.com(22740)

A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 100 gal of solution of 2% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 5% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?
:
Let x = amt of solution to be drained
and
Let x = amt of bleach to be added
:
Original 2% amt - 2% amt drained + amt of pure bleach added = 100 gal of 5%
:
.02(100) - .02x + x = .05(100)
2 - .02x + x = 5
.98x = 5 - 2
x = 3/.98
x = 3.06 gal of pure bleach should be added
:
:
Check solution:
.02(100) - .02(3.06) + 3.06 =
2 - .0612 + 3.06 = 4.9988 ~ 5